Mission Support Crew 200 Journalist Report December 6th

06-DEC-2018

Shannon told us at the beginning of the mission to notice the moment
where we felt like we were truly living on Mars. For me, I experienced
this moment. On EVA, as I followed behind my strong fellow crew
members across the snowy red hills around the MDRS, I took a deep
breath inside my spacesuit helmet. We were returning to the Hab after
a difficult hike through the hills to collect soil samples. One crew
member helped another across a small stream. As we helped each other
trek across the landscape indicating the best places to step and
lending a hand when needed, I truly felt a part of a crew on Mars.

Meanwhile, while the women of the crew went on EVA, the men were in
the Hab continuing their experiments in the lab and in the kitchen.
Lunch was fantastic (the first bread we’ve had since the mission
began). And dinner was superb! We had baked tuna melt. Although, it
was more like baked tuna crisp since we forgot to rehydrate the cheese
before putting it in the oven. Delicious all the same – we ate every
bite of the huge pan. Thanks to our Crew Astronomer (it’s been cloudy
out so now he is the Crew Chef).

Today we also had guests. We answered all their questions and showed
them around the Hab.
Oakley Jennings-Fast

Mission Support Crew 200 Sol Summary December 6th

Sol: 4

Summary Title: A productive day of science, EVAs, visitors, and outreach.

Author’s name: Lindsay Rutter

Mission Status: All crew members happy, healthy, and productive.

Sol Activity Summary: Crew ate all meals together. Science projects
continued throughout the day. Two EVAs were conducted. Crew members
welcomed more visitors to the habitat. Group photographs were taken
with crew members holding banners from girl scout troupe from Qatar.

Look Ahead Plan: No EVAs tomorrow; crew members will focus on research projects.

Anomalies in work: None reported.

Weather: Cold with snow and mud on the ground.

Crew Physical Status: No crew members reported physical problems today.

EVA: Early EVA consisted of four female crew members and was led by
female crew commander. Destination was about one mile north of the
habitat. Soil samples were collected. Later EVA consisted of three
crew members with media guests. Destination was hill next to the
habitat.

Reports to be file: N/A

Support Requested: N/A

Mission Support Crew 200 Operations Report December 6th

200 Operations Report 06-Dec-2018

SOL: 5

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 6:00 pm last night to 8:00 am. On tonight
at 6:00 pm.

Solar— SOC 82%
Diesel Reading – 25 %
Propane Reading – 40 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 390 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – not noted gallons, low.
Water (loft) – Static to Loft Pump used – yes
Water Meter: 12

Toilet tank emptied: yes

Deimos rover used: no, still not functional
Hours:
Beginning charge:
Ending charge: 74%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were not used or moved since Sunday, in charge
from Sunday afternoon.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now
parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: We had a media visit today, which we
included in our daily schedule.

Summary of internet: Nominal.

Summary of suits and radios: EVA performed. Some helmets are very
foggy (suit 1 for example), I will have to see if we can make
something for it. There is a air tube junction missing, I will have a
look tomorrow. For the cleaning spray Shannon informed us it is in the
RAM, I’ll look tomorrow.

Summary of Hab operations: Regular activities, plus Habcom support to EVA.

Summary of GreenHab operations: regular activities

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal.
Robotic Observatory – Nominal.

Summary of health and safety issues: we got back extinguishers.

Questions, concerns and requests to Mission Support:
Nothing to report

Mission Support Crew 200 EVA Report December 6th

EVA #2

Author: Makiah Eustice
Purpose of EVA: Engineering check, experiencing walking/hiking,
finding purple samples

Start time: 10:50

End time: 13:00

Narrative: Makiah performed the engineering check of the station, we
traveled until we could spot purple-like regolith. We hiked about 12
meters to collect a sample and view the landscape. We walked back.

Destination: Not-named

Coordinates (use UTM NAD27 CONUS): 518000 E – 4250000 N to 518000 E – 4252000 N

Participants: Oakley, Makiah Ilaria, Lindsay

Road(s) and routes per MDRS Map : not used, short travel

Mode of travel: walking

Astronomy Report December 6th

Astronomer: Andrew Foster

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: MDRS-WF (05 DEC)

Requested objects – Requested yesterday (05 DEC) 46/P Wirtanen (wide field),

MDRS-14 imaging update:

46/P Wirtanen (cometary coma imaging) – Active

3C 273 (Quasar) – Active

AG Peg / RW Aur (photometry measurements) – Images available (will download post-mission for processing).

Objects Viewed: None

Image attached: None.

Problems Encountered:

Poor weather conditions (Cloudy conditions continue today).

MUSK OBSERVATORY

Solar Features Observed: Short observation of Sun through eyepiece w/ high cloud conditions, no detail visible. Carried out short test of Imaging Source DBK camera (no image captured).

Notes

Carried out minor repair on Solar observatory dome – loose seal secured (see attached photos).

Both observatories are clear of snow this evening.

End report.

Greenhab Report December 6th

Crew 200 Greenhab Report 06-DEC-2018

Greenhab Officer:
Makiah Eustice

Environmental Control:
Heating
40% Shade Cloth on

Average Temperatures:
Low: 14 C
High: 40 C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
2.5 gallons

Water in Blue Tank (lbs):
29%

Times of Watering for Crops:
1200

Changes to crops:
Took out wilting snap peas.

Narrative:
Soaking Snap peas, salad green mix, and sunflower seeds to sow
tomorrow. Got some Oregano for our dinner and planning on a good salad
tomorrow!

Harvest:
5g Oregano

Support/Supplies Needed:
Let me know if there is any comment on a pulley system that could be
implemented to easily lower and raise the lights.